OpenStudy (anonymous):

if the net price of a stove is $900 and the trade discount raye is 40% the list price is....?

OpenStudy (anonymous):

Okay not sure if I understand but if the net price is $900 and you received 40% discount off the list price then, 900/x=60/100 which is 60x=90000, x=$1,500... which is the original list price...

OpenStudy (anonymous):

Hey Gandalfwiz

OpenStudy (anonymous):

the single equivelent discount rate of the trade discount 5/4/1 is...?